# Stable electron-irradiated [1-13C]alanine radicals for metabolic imaging with Dynamic Nuclear Polarisation Short Title: Electron-irradiated radicals for DNP

### v2

@@ 4.4 In Vivo Experiments
Two healthy male Sprague-Dawley rats were injected with one shot of each type of hyperpolarised alanine sample via a tail vein catheter. The time between each injection was approximately two hours. The respiratory rate and body temperature of each rat were monitored using an MRI compatible small-animal monitoring system (Small Animal Instruments Inc, USA) whilst the rats were under anaesthesia with 2.5-3% sevoflurane in 1 l/min medical air. Both rats were scanned using a dual-tuned 13C/1H volume rat coil at 3 T (GE MR750, GE Healthcare). One rat was scanned using a slice-selective spectroscopic acquisition, the other imaged using a 2D MR Spectroscopic Imaging (MRSI) sequence, EPSI, using fidall. (?)
Prior to these scans, a urea phantom was used to calibrate the applied flip angles, and proton images were acquired for anatomical reference. For the slice-selective spectroscopic acquisition, 80 mm slice thickness was used to ensure the whole abdomen of the rat was imaged. 1 ml of hyperpolarised sample was injected over approximately 5 s, and the scan started upon the start of the injection. A 30  flip angle was applied with a repetition time (TR) of 3 s for a total of 5 minutes (100 time points). For the MRSI, 1 ml of hyperpolarised substrate was injected over approximately 5 s, and the scan started upon the start of the injection. A $16\times 16$ matrix was chosen, and a 15  flip angle was applied with a TR of 64 ms for a total of 2 minutes (8 time points and approximately 16 s per time point). The pH of the injected solutions was measured to ensure physiological compatibility, with values ranging from 6.7 to 7.4, well within the acceptable range for intravenous administration. All animal experiments were conducted following appropriate independent ethical review and under appropriate licensing regimes in Denmark.
@@ 4.5 Electron Paramagnetic Resonance
An X-band continuous wave (CW) spectrometer (EMXmicro, Bruker BioSpin; Super High Sensitivity rectangular cavity probe head) was used to record EPR spectra at room temperature. For recording spectra at 5 K and measurement of $T_{1e}$ and $T_{2}$, an X-band pulsed spectrometer (E680, Bruker BioSpin) was used. The Chromium(III) impurity in Magnesium Oxide powder, a common standard in EPR, (?, ?) was used as a frequency reference. Further details are provided in SI §6.7.
@@ 4.6 X-ray Diffraction
XRD data were compared to a previously reported crystal structure of alanine, obtained via neutron scattering (?), and a Rietveld refinement was undertaken in Maud using a spherical harmonic texture basis (?). Further details are provided in the SI §6.9. [7890 Words]
@@ 5 Figures and Tables
Figure 1: Radical generation and quantification. (A) An ultra-high dose rate optimised in-house developed linear accelerator was used to bombard (B) a target containing polycrystalline [1-13C]alanine with 6 MeV electrons. (C) These generate Stable Alanine Radicals (SAR) detectable by EPR in a dose-dependent fashion compared to (D) TEMPO as a concentration standard. (E) These EPR data were acquired in a single CW EPR scan at 9.38611 GHz with a modulation frequency of 100 kHz; and sweep rate of 0.816 mT/s. The 13C-hyperfine interaction in the SARs, significantly alters the EPR spectrum compared to natural abundance (NA) alanine, which we modelled (F) via simulations in EasySpin, shown frequency-corrected to a standard frequency of 9.6 GHz, with our 13C-model in black (detailed in SI §6.8).
@@
Figure 2: DNP and Glycerol addition. (A) DNP build-up profiles of 13C-NMR signal arising from different high dose rate irradiated samples at 3.35 T; we found that dispersing irradiated polycrystalline alanine powder in glycerol produced substantially higher polarisation. (B) Longitudinal measurements of the spin counts showed that there was no significant effect on the endogenous radical concentration with increased duration of storage in anhydrous glycerol up to 4 months, in sharp contrast to pyruvic acid/trityl radical mixtures which quench within hours (mean $\pm$ SD of three technical replicates). (C) Alanine/glycerol mixtures did not significantly alter acquired EPR data at room temperature (RT), but at (D) 5 K the electronic environment altered considerably, and the EPR spectrum recorded shows significant changes from RT data. These data were obtained via either CW EPR as previously (RT) at 9.38611 GHz or at 5 K on a pulsed instrument at 9.76344 GHz and have been frequency-corrected to 9.6 GHz in (D) or 9.4 GHz in (C), for display on a common scale; c.f. SI §6.7.
@@
Figure 3: DNP at 6.7 T and partial order(A) The frequency sweep profile of DNP with alanine/glycerol mixtures showed qualitative differences with increasing radiation dose, which we hypothesise is due to cooperativity. (B) Polarisation build-up as a function of time showed profound biexponential behaviour under these conditions, distinct from narrow linewidth radicals such as trityl; all arising from 13C-NMR on [1-13C]alanine. (C, Di) X-ray diffractometery measurements of both samples were able to demonstrate that polycrystalline alanine remained dispersed (not dissolved) in glycerol and was well described by (Dii) existing crystal structures with slight refinement but had profound crystallographic texture, indicative of an ordered crystalline arrangement within the glycerol matrix. (E) The efficiency of DNP under these conditions is surprising, comparable to narrow linewidth trityl radicals. Further experimental details are provided in the SI.
@@
Figure 4: Polarisation optimisation and high-field comparisons with trityl. (A) At 6.7 T, a unimodal, always-positive frequency sweep curve was observed. Pulsed modulation with a swept-frequency microwave pulse as illustrated obtained a significant increase in nuclear polarisation. (B) The molecular orientation of the alanine radical matters significantly for the spectral density function $g(\omega)$ as defined by Wenckebach, (?) simulated here using our model derived from powder 13C alanine EPR data and assuming a partial ordering parameter $\lambda$. (C) The resultant limiting polarisation obtained was high, and comparable to the use of trityl radicals with alanine, with biexponential fit shown. (D) After dissolution, spontaneous quenching of endogeneous radicals formed by electron irradiation resulted in a lengthened $T_{1}$ (77 s) compared to that with trityl (40 s) at 1.4 T (fits shown in C). Further experimental details are provided in the SI.
@@
Figure 5: In vivo demonstration in the rat. (A, B) Stacked and (C, D) summed spectra from slice-selective spectroscopic acquisitions following separate tail-vein injections of [1-13C]alanine polarised via 38 mM OX063 trityl radical and hyperpolarised irradiated 13C-alanine. (C, D) Inset: a magnification of the spectra presented in the panels to highlight the intensities and positions of the downstream metabolites, whose chemical shift values match those reported in the literature for 13C-lactate and 13C-pyruvate. (E, F) After quantification by AMARES, metabolite dynamics were comparable between the two methods of preparation. (G) MR Spectroscopic Imaging (MRSI) images obtained in a separate rat were likewise near identical, validating the use of the technique for in vivo imaging of redox potential, alanine images of which are shown concatenated over time. Further experimental details are provided in the SI.
Table 1: Summary of DNP build-up parameters from biexponential fits across different experimental conditions and radical sources. Build-up profiles follow the biexponential model: $y=P[A(1-e^{-t/T})+(1-A)(1-e^{-t/\tau})]$, where $P$ is the build-up constant, AA is the fast component weighting, $T$ is the fast component time constant, and $\tau$ is the slow component time constant. It was not possible to fit 3.35 T data to a biexponential curve stably; and likewise a biexponential fit to trityl data at 6.7 T returned monoexponential behaviour. Arbitrary units are comparable within experiments but not between them.
Fig. Sample/Condition Dose Build-up Fast Slow
(kGy) Amplitude Time $T$ Time $\tau$ Weighting AA
$P$ (a.u.) (s) (h)
2a 10 kGy (glycerol plug) 10 $9.59\times 10^{2}\dagger$ Monoexponential fit poor; $\tau_{\text{mono}}=3171342310$ s
100 kGy (glycerol plug) 100 $6.28\times 10^{3}\dagger$ ($\tau_{\text{mono}}=19236$ s)
100 kGy (glycerol mix) 100 $2.37\times 10^{4}\dagger$ ($\tau_{\text{mono}}=2836$ s)
3b 30 kGy (glycerol mix) 30 $1.52\times 10^{6}$ $22.9$ $0.92$ $0.264$
50 kGy (glycerol mix) 50 $8\times 10^{6} 22.22$2.2 $2.69$ $0.193$
70 kGy (glycerol mix) 70 $5.11\times 10^{6}$ $14.9$ $6.88$ $0.099$
[…]
70 kGy (glycerol mix) 70 $1.7\times 10^{1}$ 2020 $0.481$ $0.088$
Notes: $\dagger$ These data are at 3.35 T.
$\ddagger$ OX063 trityl radical shows purely monoexponential behaviour ($A=0$ in one fit; identical time constants in the other).
@@ Acknowledgments
We would like to acknowledge support from the Novo Nordisk Foundation (NNF21OC0068683); the British Heart Foundation (refs. RG/11/9/28921, RE/13/1/30181, FS/19/18/34252, and FS/14/17/30634); the Wellcome Trust (221805/Z/20/Z); the Engineering and Physical Sciences Research Council (EPSRC) and Medical Research Council (MRC) (EP/L016052/1); and the Colleges of St. Hugh’s and Somerville in the University of Oxford. We would furthermore like to acknowledge the University of Oxford British Heart Foundation Centre for Research Excellence (RE/13/1/30181) and the NHS National Institute for Health Research Oxford Biomedical Research Centre programme. The views expressed are those of the authors and not necessarily those of the NIHR or the Department of Health and Social Care. AMB is grateful to The Royal Society and the EPSRC for a Dorothy Hodgkin Fellowship (DH160004 and DHF/R/221018), and the University of Manchester for a Dame Kathleen Ollerenshaw Fellowship. AMB also thanks the Royal Society of Chemistry for a Community for Analytical and Measurement Science fellowship (CAMS Fellowship 2020 ACTF ref 600310/09). AMB is grateful for support from National Research Facility for EPR (EP/W014521/1, NS/A000055/1, EP/V035231/1, and EP/S033181/1). We also gratefully acknowledge Cambridge Isotope Laboratories, Inc, who provided a gift of [1-13C]alanine to JJM and Z.R. We also thank Christiane Timmel for useful discussions about quantitative EPR spectroscopy. JJM would additionally like to thank Professor Elspeth Garman and Professor Ian Carmichael for both dinner and many useful discussions about radiation chemistry in crystals. The practical assistance of the University of Oxford Department of Physics Mechanical Workshops is also kindly noted.
@@
CHER and JYCL conceptualised research, conducted experiments, analysed data, and drafted the manuscript. ESSH, NVC, and DAD assisted with sample preparation and in vivo/high field DNP experiments; BWC and SS assisted with low-field DNP experiments. KP, IT, BV, CHER undertook electron linear accelerator experiments for sample irradiation, including design and manufacture of equipment. JL, CHER and JJM undertook X-ray diffraction experiments and their interpretation. WM, and AMB provided expertise in EPR spectroscopy and data interpretation. ZR assisted with modelling. LBB and CL supervised the in vivo experiments and contributed to experimental design. JYCL, DJT and JJM conceived and supervised the project, secured funding, analysed data, and finalised the manuscript. All authors reviewed, edited and approved the final manuscript.
@@
JYCL, CL, JJM, CHER and DT are listed as authors on a filed patent pending examination (EP24168322.6) pertaining to the use of ultra-high dose rate electron irradiation for preparing hyperpolarised contrast agents. All other authors have no conflicts of interest to declare.
@@
Crystallographic data pertaining to this project are enclosed as CIF files as SI to this paper, as is EPR raw data. Materials (i.e. ultra-high dose rate irradiated labelled metabolites) are available to others upon the signing of a collaboration and material transfer agreement. All other data are available in the main text or the supplementary materials.
@@
Tables S1 to S2
References (123-0)
Data S1 to S2
@@
Stable electron-irradiated [1-13C]alanine radicals for metabolic imaging with Dynamic Nuclear Polarisation
Catriona H. E. Rooney, Justin Y. C. Lau, Esben S. S. Hansen, Nichlas Vous Christensen, Duy A. Dang, Kristoffer Petersson, Iain D. C. Tullis, Borivoj Vojnovic, Sean Smart, William Myers, Zoe Richardson, Jarrod Lewis, Brett W. C. Kennedy, Alice M. Bowen, Lotte Bonde Bertelsen, Christoffer Laustsen, Damian J. Tyler, and Jack J. Miller∗‡
Corresponding author. Email: jack.miller@physics.org $^{\dagger},\ddaggerThese$ authors contributed equally to this work.
@@ 6.1 Irradiation
Renderings of the sample holder used to hold the sample during the irradiation are shown Fig. S1A. To produce the samples, a block of aluminium with two sample spaces was loaded with dry alanine powder. The powder was secured in place by taping two ordinary glass microscope slides either side of the aluminium block. The glass-aluminium-glass sandwich was held in an upright position by microscope slide spring clips within another aluminium component that itself was connected to a high precision linear slider. The integration of the slider and two stoppers into the design facilitated the centring of each sample space within the beam path. To increase the homogeneity of the irradiation exposure across the total volume of each sample space, the glass-aluminium-glass sandwich was flipped 180° after half of the total desired irradiation dose had been delivered. Irradiation doses over the range 10–100 kGy were applied to produce samples with a range of different endogenous radical concentrations.
A radiation transport simulation was undertaken in TOPAS, (?, ?) an open-source radiation transport code optimised for medical physics applications based on Geant4, written by CERN. (?) As shown in Fig. S1B, C, this was used to optimise dose distribution across the sample with regards to the ‘pile up’ dose deposited either in the glass coverslips or aluminium target; the preferred deposition function is that of a ‘top hat’ within the sample but it was found that appropriate spacing and modulation of design produced a reasonable distribution function (Fig. S1D). A significant amount of charge transfer is expected to take place within the electron bombardment process. In order to obtain a more homogeneous dose distribution within the sample, we rotated the sample by 180  in the middle of the irradiation process; after irradiation, the sample was mixed by hand, further homogenising it.
Figure S1: Details of the overall process of e-beam irradiation. A: A sample holder for the irradiation of alanine samples using a 6 MeV high-dose rate electron linear accelerator was designed. Renderings of the sample holder comprising (1) an aluminium endplate with a 5 cm diameter aperture for the electron beam, (2) a rectangular block of aluminium — with two cylindrical holes each acting as a sample space — interposed between two glass microscope slides and secured upright with microscope slide spring clips to (3) an L-shaped aluminium block attached onto (4) a high precision linear slider. (5) Two stoppers were positioned either side of the slider to facilitate the positioning of each sample. B, C: Simulated Monte-Carlo radiation transport characteristics of the designed target demonstrating appropriate dose deposition inside the compressed alanine pellet (red denotes leptons; green photons); together with D: predicted dose deposition as a function of depth, normalised such that the total dose is unity.

### v3

@@ 4.4 In Vivo Experiments
Two healthy male Sprague-Dawley rats were injected with one shot of each type of hyperpolarised alanine sample via a tail vein catheter. The time between each injection was approximately two hours. The respiratory rate and body temperature of each rat were monitored using an MRI compatible small-animal monitoring system (Small Animal Instruments Inc, USA) whilst the rats were under anaesthesia with 2.5-3% sevoflurane in 1 l/min medical air. Both rats were scanned using a dual-tuned 13C/1H volume rat coil at 3 T (GE MR750, GE Healthcare). One rat was scanned using a slice-selective spectroscopic acquisition, the other imaged using a 2D MR Spectroscopic Imaging (MRSI) sequence, EPSI, using fidall.
Prior to these scans, a urea phantom was used to calibrate the applied flip angles, and proton images were acquired for anatomical reference. For the slice-selective spectroscopic acquisition, 80 mm slice thickness was used to ensure the whole abdomen of the rat was imaged. 1 ml of hyperpolarised sample was injected over approximately 5 s, and the scan started upon the start of the injection. A 30  flip angle was applied with a repetition time (TR) of 3 s for a total of 5 minutes (100 time points). For the MRSI, 1 ml of hyperpolarised substrate was injected over approximately 5 s, and the scan started upon the start of the injection. A $16\times 16$ matrix was chosen, and a 15  FA was applied with a TR of 64 ms for a total of 2 minutes (8 time points and approximately 16 s per time point). The pH of the injected solutions was measured to ensure physiological compatibility, with values ranging from 6.7 to 7.4, well within the acceptable range for intravenous administration. All animal experiments were conducted following appropriate independent ethical review and under appropriate licensing regimes in Denmark.
@@ 4.5 Electron Paramagnetic Resonance
An X-band continuous wave (CW) spectrometer (EMXmicro, Bruker BioSpin; Super High Sensitivty rectangular cavity probe head) was used to record EPR spectra at room temperature. For recording spectra at 5 K and an estimate of $T_{1e}$, an X-band CW spectrometer (E680, Bruker BioSpin) was used. Magnesium Oxide powder, a common standard in EPR, (?) was used as a frequency reference.
@@ 4.6 X-ray Diffraction
XRD data were compared to a previously reported crystal structure of alanine, obtained via neutron scattering (?), and a Rietveld refinement was undertaken in Maud using a spherical harmonic texture basis (?). Further details are provided in the SI §6.6.
[6167 Words]
@@ 5 Figures and Tables
Figure 1: Radical generation. (A) An ultra-high dose rate optimised in-house developed linear accelerator was used to bombard (B) a target containing polycrystalline [1-13C]alanine with 6 MeV electrons. (C) These generate Stable Alanine Radicals (SAR) detectable by EPR in a dose-dependent fashion compared to (D) TEMPO as a concentration standard. (E) The 13C-hyperfine interaction is strong for the SARs, significantly altering the EPR spectrum compared to natural abundance (NA) alanine, which we quantified (F) via simulations in EasySpin.
@@
Figure 2: (A) DNP build-up profiles of different high dose rate irradiated samples at 3.35 T; we found that dispersing irradiated polycrystalline alanine powder in glycerol produced substantially higher polarisation. (B) Longitudinal measurements of the spin counts showed that there was no significant effect on the endogenous radical concentration with increased duration of storage in anhydrous glycerol up to 4 months, in sharp contrast to pyruvic acid/trityl radical mixtures which quench within hours (mean $\pm$ SD of three technical replicates). (C) Alanine/glycerol mixtures did not significantly alter acquired EPR data, but at (D) 5 K the electronic environment altered considerably
@@
Figure 3: (A) The frequency sweep profile of DNP with alanine/glycerol mixtures showed qualitative differences with increasing radiation dose, which we hypothesise is due to cooperativity. (B) Polarisation build-up as a function of time showed profound biexponential behaviour under these conditions, distinct from narrow linewidth radicals such as trityl. (C, Di) X-ray diffractometery measurements of both samples were able to demonstrate that polycrystalline alanine remained dispersed (not dissolved) in glycerol and was well described by (Dii) existing crystal structures with slight refinement but had profound crystallographic texture, indicative of an ordered crystalline arrangement within the glycerol matrix. (E) The efficiency of DNP under these conditions is surprising, comparable to narrow linewidth trityl radicals.
@@
Figure 4: (A) At 6.7 T, a unimodal, always-positive frequency sweep curve was observed. Pulsed modulation with a swept-frequency microwave pulse as illustrated obtained a significant increase in nuclear polarisation. (B) The molecular orientation of the alanine radical matters significantly for the spectral density function $g(\omega)$ as defined by Wenckebach, (?) simulated here using parameters from powder 13C alanine EPR data and assuming a partial ordering parameter $\lambda$. (C) The resultant limiting polarisation obtained was high, and comparable to the use of trityl radicals with alanine. (D) After dissolution, spontaneous quenching of endogeneous radicals formed by electron irradiation resulted in a lengthened $T_{1}$ (77 s) compared to that with trityl (40 s) at 1.4 T.
@@
Figure 5: (A, B) Stacked and (C, D) summed spectra from slice-selective spectroscopic acquisitions following separate tail-vein injections of [1-13C]alanine polarised via 38 mM OX063 trityl radical and hyperpolarised irradiated 13C-alanine. (C, D) Inset: a magnification of the spectra presented in the panels to highlight the intensities and positions of the downstream metabolites, whose chemical shift values match those reported in the literature for 13C-lactate and 13C-pyruvate. (E, F) After quantification by AMARES, metabolite dynamics were comparable between the two methods of preparation. (G) MR Spectroscopic Imaging (MRSI) images obtained in a separate rat were likewise near identical, validating the use of the technique for in vivo imaging of redox potential, alanine images of which are shown concatenated over time.
@@ Acknowledgments
We would like to acknowledge support from the Novo Nordisk Foundation (NNF21OC0068683); the British Heart Foundation (refs. RG/11/9/28921, RE/13/1/30181, FS/19/18/34252, and FS/14/17/30634); the Wellcome Trust (221805/Z/20/Z); the Engineering and Physical Sciences Research Council (EPSRC) and Medical Research Council (MRC) (EP/L016052/1); and the Colleges of St. Hugh’s and Somerville in the University of Oxford. We would furthermore like to acknowledge the University of Oxford British Heart Foundation Centre for Research Excellence (RE/13/1/30181) and the NHS National Institute for Health Research Oxford Biomedical Research Centre programme.
The views expressed are those of the authors and not necessarily those of the NIHR or the Department of Health and Social Care. We also gratefully acknowledge Cambridge Isotope Laboratories, Inc, who provided a gift of [1-13C]alanine to J.J.M. and Z.R. We also thank Christiane Timmel for useful discussions about quantitative EPR spectroscopy. JJM would additionally like to thank Professor Elspeth Garman and Professor Ian Carmichael for both dinner and many useful discussions about radiation chemistry in crystals. The practical assistance of the Department of Physics Mechanical Workshops is also kindly noted.
@@
C.H.E.R. and J.Y.C.L. conceptualised research, conducted experiments, analysed data, and drafted the manuscript. E.S.S.H., N.V.C., and D.A.D. assisted with sample preparation and in vivo/high field DNP experiments; B.W.C and S.S. assisted with low-field DNP experiments. K.P., I.T., B.V., C.H.E.R. undertook electron linear accelerator experiments for sample irradiation, including design and manufacture of equipment. J.L, C.H.E.R. and J.J.M undertook X-ray diffraction experiments and their interpretation. W.M., and A.B. provided expertise in EPR spectroscopy and assisted with data interpretation. Z.R. assisted with modelling. L.B.B. and C.L. supervised the in vivo experiments and contributed to experimental design. J.Y.C.L, D.J.T. and J.J.M. conceived and supervised the project, secured funding, analysed data, and finalized the manuscript. All authors reviewed, edited and approved the final manuscript.
@@
JYCL, CL, JJM, CHER and DT are listed as authors on a filed patent (EP24168322.6) pertaining to the use of ultra-high dose rate electron irradiation for preparing hyperpolarised contrast agents. All other authors have no conflicts of interest to declare.
@@
Crystallographic data pertaining to this project are enclosed as CIF files as SI to this paper. Materials (i.e. ultra-high dose rate irradiated labelled metabolites) are available to others upon the signing of a collaboration and material transfer agreement. All other data are available in the main text or the supplementary materials.
@@
Table S1
References (7-0)
Data S1
@@
Stable electron-irradiated [1-13C]alanine radicals for clinically viable metabolic imaging with Dynamic Nuclear Polarization
Catriona H. E. Rooney et al. Corresponding author. Email: jack.miller@physics.org
@@ 6.1 Irradiation
Renderings of the sample holder used to hold the sample during the irradiation are shown in panel (a) of Fig. S1. To produce the samples, a block of aluminium with two sample spaces was loaded with dry alanine powder. The powder was secured in place by taping two ordinary glass microscope slides either side of the aluminium block. The glass-aluminium-glass sandwich was held in an upright position by microscope slide spring clips within another aluminium component that itself was connected to a high precision linear slider. The integration of the slider and two stoppers into the design facilitated the centring of each sample space within the beam path. To increase the homogeneity of the irradiation exposure across the total volume of each sample space, the glass-aluminium-glass sandwich was flipped 180° after half of the total desired irradiation dose had been delivered. Irradiation doses over the range 10–100 kGy were applied to produce samples with a range of different endogenous radical concentrations.
A radiation transport simulation was undertaken in TOPAS, (?, ?) an open-source radiation transport code optimised for medical physics applications based on Geant4, written by CERN. (?) As shown in Fig. S1b, c, this was used to optimise dose distribution across the sample with regards to the ‘pile up’ dose deposited either in the glass coverslips or aluminium target; the preferred deposition function is that of a ‘top hat’ within the sample but it was found that appropriate spacing and modulation of design produced a reasonable distribution function (Fig. S1d). In order to obtain a more homogeneous dose distribution within the sample, we rotated the sample by 180  in the middle of the irradiation process; after irradiation, the sample was mixed by hand, further homogenising it.
Figure S1: a: A sample holder for the irradiation of alanine samples using a 6 MeV high-dose rate electron linear accelerator was designed. a Renderings of the sample holder comprising (1) an aluminium endplate with a 5 cm diameter aperture for the electron beam, (2) a rectangular block of aluminium — with two cylindrical holes each acting as a sample space — interposed between two glass microscope slides and secured upright with microscope slide spring clips to (3) an L-shaped aluminium block attached onto (4) a high precision linear slider. (5) Two stoppers were positioned either side of the slider to facilitate the positioning of each sample. b, c: Simulated Monte-Carlo radiation transport characteristics of the designed target demonstrating appropriate dose deposition inside the compressed alanine pellet (red denotes leptons; green photons); together with (d) predicted dose deposition as a function of depth, normalised such that the total dose is unity.